### v2

# TempoControl: Temporal Attention Guidance for Text-to-Video Models

Abstract
Recent advances in generative video models have enabled the creation of high-quality videos based on natural language prompts. However, these models frequently lack fine-grained temporal control, meaning they do not allow users to specify when particular visual elements should appear within a generated sequence. In this work, we introduce TempoControl, a method that allows for temporal alignment of visual concepts during inference, without requiring retraining or additional supervision. TempoControl utilizes cross-attention maps, a key component of text-to-video diffusion models, to guide the timing of concepts through a novel optimization approach. Our method steers attention using three complementary principles: aligning its temporal pattern with a control signal (correlation), adjusting its strength where visibility is required (magnitude), and preserving semantic consistency (entropy). TempoControl provides precise temporal control while maintaining high video quality and diversity. We demonstrate its effectiveness across various applications, including temporal reordering of single and multiple objects, action timing, and audio-aligned video generation. Please see our project page for more details: https://shira-schiber.github.io/TempoControl/.
Figure 1: Applications of our inference-time temporal control method (TempoControl), which enables multi-object and single-object control, motion control, and audio alignment. Words that are controlled to appear throughout the entire video are shown in blue. Words that appear according to a temporally varying control signal are shown in orange.
1 Introduction
Generative video models have improved dramatically, enabling the synthesis of high-quality, temporally coherent videos from natural language descriptions [22, 6, 34, 3, 23]. These models have enabled creative applications in animation, design, and virtual content creation by generating visually compelling sequences that align with textual descriptions. However, while current models offer strong spatial consistency and global semantics, they lack fine-grained temporal control. Specifically, guiding the model to generate visual elements at specific time points is particularly challenging. This includes making an object appear in the middle of a scene or aligning visual aspects of lighting with the sound of thunder (see Figure 1).
Recent works on generative video modeling have proposed various spatial and motion control mechanisms, including object masking, layout conditioning, or added camera controls [10, 25, 9, 11]. In contrast, temporal control has received significantly less attention. Incorporating such control typically requires additional supervision in the form of temporally annotated datasets. Since video data is often scarce, extending annotations to capture temporal variation is expensive and usually impractical at scale. Synthetic data could be used; however, generating diverse and temporally accurate video-text pairs, especially for abstract concepts such as movement, is a highly non-trivial challenge.
In this work, we propose a lightweight yet effective inference-time control method, TempoControl, which introduces temporal control for generated video. Instead of retraining the generative model, TempoControl leverages a property already embedded in these systems: the text-vision cross-attention mechanism.
Cross-attention layers inherently encode strong signals about which specific words in the prompt are realized in each generated frame. By directly influencing these attention maps during the denoising process, we can shift the temporal alignment of visual concepts without sacrificing quality or requiring additional data. The approach is data-efficient and preserves both the fidelity and diversity of the original generation process.
Attention control is achieved by carefully steering the latent variables during the diffusion process. At each denoising step, we apply a few stochastic gradient descent iterations [7, 5] until a satisfactory level of temporal alignment is achieved, notably, without updating model parameters.
To guide temporal alignment, TempoControl introduces a novel loss with three complementary components, based on Pearson correlation, magnitude, and entropy. The Pearson term encourages the temporal shape of attention to match that of an external control signal, operating on the normalized attention vector to align the presence of a concept with the desired timing. However, since correlation is scale-invariant, it may yield high scores even when attention values are too low to render the object visible. To mitigate this, we incorporate a magnitude term that directly promotes stronger attention in frames where the temporal signal is high, and suppresses it elsewhere. Additionally, we introduce an entropy regularization term to maintain spatial focus, ensuring that when attention is activated, it remains coherent and not diffusely spread across the frame.
Beyond single-object scenarios, as shown in Figure 1, our method generalizes well to multi-object scenes. In such cases, it is crucial to maintain the visual grounding of non-target entities. We find that preserving one object (e.g., the bird) while selectively manipulating another (e.g., the cat) helps retain the overall scene structure. In addition to nouns, our method can also manipulate verbs (e.g., clapping), effectively modifying the motion depicted in the video. Finally, we explore the potential of aligning external audio cues with video generation. While we focus on simple cases involving a single audio event (e.g., a lightning strike), we observe promising results, suggesting that our approach may naturally extend to zero-shot multimodal control without additional training.
Our contributions are threefold: (i) We introduce the first inference-time method for temporally controlling the appearance of specific words in text-to-video diffusion models, without additional supervision or model finetuning, allowing temporal control while preserving the fidelity and diversity of existing models. (ii) We propose a set of spatiotemporal attention losses that shape both the timing and strength of word grounding, together with an entropy-based regularization term that maintains semantic consistency. (iii) We demonstrate broad applicability across multi-object control, motion timing, and audio-to-video alignment, and introduce a new metric for temporal alignment, a dimension not covered by current benchmarks.
Figure 2: Motivation for our approach. Top: We show a video generated for the prompt “The video begins with a serene view centered on the cat, with no sign of the dog. In the second half, the dog unexpectedly appears, altering the dynamic of the scene.” The top row displays the attention maps for the tokens cat and dog, extracted from the denoising step t=3t{=}3, for frames j=2,6,10,14,19j{=}2,6,10,14,19. On the left, the video is generated without our optimization. Despite the prompt specifying that the dog should appear in the second half, it appears early. This behavior is common, as Wan 2.1 often fails to depict objects or movements according to temporal cues in the prompt. On the right, after applying our conditioning method, the dog correctly appears in the second half of the video. Bottom: Temporal attention $a_{i,j}^{t}$ (blue) versus target mask $m_{i,j}$ (orange), with the corresponding Pearson correlation loss shown.
2 Related Work
Attention-based Control. Attention manipulation has become a powerful tool for controlling semantics in text-to-image generation. Methods such as Attend-and-Excite [7, 17, 4, 8] perform inference-time optimization over cross-attention maps to ensure that all textual entities are faithfully represented in the generated image. Stable Flow [1] further introduces a training-free editing framework that selectively injects attention features into key layers of a diffusion transformer. Additional approaches, including Prompt-to-Prompt [12], leverage attention restructuring for fine-grained, text-driven edits.
Inference time optimization. Inference-time optimization for text-to-video diffusion has only recently attracted attention. [16] pre-compute a spatio-temporal sketch and inject plan-consistent noise during sampling to tighten prompt alignment. [28] converts a single keyframe into a textured, deformable 3D mesh that is re-rendered at every denoising step to maintain consistency in the garment’s appearance. [21] derive variance-based optical-flow cues from the model’s latent predictions and add a flow-coherence loss on-the-fly, smoothing motion without retraining. [18] let a large language model generate scene graphs and temporal layouts that act as attention masks, improving complex spatio-temporal grounding. While these methods improve fidelity and consistency without necessitating additional training, none offer explicit, user-defined control over the timing of textual element unfolding.
Controllable Video Generation Controllable video generation has advanced rapidly, with methods exploring a broad set of control modalities. Early multimodal systems, such as VideoComposer [26], incorporate depth, sketch, and motion vector guidance to steer video synthesis. More recent approaches introduce explicit camera and object motion control on top of text-to-video generation [27, 11, 15, 33, 9, 2]. Another line of work investigates the controllability of image-to-video object motion [25, 32]. While effective, these training approaches rely on large paired datasets and extensive supervision, which can be costly and limit generalization.
A parallel line of work studies training-free controllability, including CamTrol [13] for camera motion, MotionClone and DiTFlow [19, 20] for transferring motion from reference videos, and MOFT [31] for editing motion at inference. These rely on dense motion trajectories, but cannot specify when a concept should appear or disappear.
The most related work, MinT [30], arranges multiple textual instructions in temporal order by learning when each should begin. However, MinT provides only coarse event sequencing, whereas our method offers fine-grained timing for individual concepts. Moreover, MinT requires fine-tuning on temporally annotated caption datasets centered on human activities, which limits its generality across diverse prompts. Our approach operates entirely at inference time and generalizes across settings without any training or domain-specific supervision.
3 Preliminaries
We build TempoControl on top of Wan 2.1 [23], a state-of-the-art diffusion model that serves as our backbone for text-to-video generation. Let $v\in\mathbb{R}^{T_{v}\times H_{v}\times W_{v}\times 3}$ represent the input video, where $T_{v}$ is the number of frames, $H_{v}\times W_{v}$ denotes the spatial resolution of each frame, and the final dimension corresponds to the three RGB color channels. A 3D causal variational autoencoder first encodes the video into a latent tensor
$z\in\mathbb{R}^{T^{\prime}\times H^{\prime}\times W^{\prime}\times C},$ (1)
where $T^{\prime}$, $H^{\prime}$, and $W^{\prime}$ denote the temporal and spatial dimensions of the latent representation, which are reduced relative to the input video. Flattening the spatial and temporal dimensions, $z$ is reshaped into a sequence of $n_{v}$ video tokens, where $n_{v}=T^{\prime}\cdot H^{\prime}\cdot W^{\prime}$, and each token is a $C$-dimensional feature vector representing a specific spatiotemporal region of the video.
The text prompt $\mathcal{P}=[p_{1},\dots,p_{n_{p}}]$ is embedded using a pretrained text encoder. For brevity, we refer to tokens as words, although a word may consist of multiple subtokens. All operations described for words are applied to the average of their corresponding subtokens. Video synthesis is then performed as a denoising diffusion process in the latent space, where the model iteratively refines latent variables $\{z_{t}\}_{t=T}^{0}$, gradually denoising $z_{T}$ into $z_{0}$. At each timestep $t$, a transformer-based network $\operatorname{DiT}(z_{t},\mathcal{P},t)$ predicts the noise in $z_{t}$, conditioned on the prompt $\mathcal{P}$ and the timestep $t$.
3.1 Cross-attention in video diffusion models
A key challenge in text-to-video generation is maintaining semantic and temporal alignment with the input prompt. In diffusion-based architectures, this alignment is achieved through cross-attention layers within the transformer blocks at each denoising step $t$.
During cross-attention, video tokens act as queries that attend to text tokens, which serve as keys and values. This mechanism enables the model to inject contextually relevant linguistic information into each video token representation. The attention weights are computed via scaled dot products between queries and keys, producing a set of attention tensors across layers, $A^{t}\in\mathbb{R}^{L\times h\times n_{v}\times n_{p}}$, where $L$ is the number of cross-attention layers, $h$ is the number of attention heads, $n_{v}$ is the number of video tokens, and $n_{p}$ is the number of text tokens. A softmax is applied along the text-token dimension, yielding a probability distribution over prompt words for each video token, per head and per layer.
To obtain a global view of token-level influence, we average attention maps across heads and cross-attention layers, producing an aggregate attention matrix $\bar{A}^{t}\in\mathbb{R}^{n_{v}\times n_{p}}$. This matrix quantifies the influence of each text token on each video token.
The effect of a specific word $i$ is measured by extracting the corresponding column $\bar{A}^{t}_{i}\in\mathbb{R}^{n_{v}}$. The vector $\bar{A}^{t}_{i}$ can be reshaped at temporal index $j$, yielding $\bar{A}^{t}_{j,i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$, which represents the spatial influence of word $i$ at temporal position $j$ and denoising step $t$. Figure 2 illustrates this for the words cat and dog in the prompt “A cat and a dog. In the second half of the video, the dog appears.” Attention maps are upsampled to frame resolution for display.
The attention map captures the spatiotemporal grounding of words in the generated video; TempoControl is specifically designed to steer the temporal influence of a token. Thus, we focus on the scalar attention value $\hat{A}^{t}_{j,i}=\langle\bar{A}^{t}_{j,i}\rangle_{x,y}\in\mathbb{R}$, where $\langle\cdot\rangle_{x,y}$ denotes summation over the spatial coordinates $x\in[1,H^{\prime}]$ and $y\in[1,W^{\prime}]$. Intuitively, $\hat{A}^{t}_{j,i}$ measures the influence of word $i$ on the latent representation at temporal position $j$ during denoising step $t$.
To aggregate the temporal attention pattern of word $i$ at denoising step $t$, we collect the scalar values across frames into a vector:
$a_{i}^{t}=[\hat{A}^{t}_{1,i},\hat{A}^{t}_{2,i},\ldots,\hat{A}^{t}_{T^{\prime},i}]\in\mathbb{R}^{T^{\prime}}.$ (2)
We omit the first frame due to unstable attention values that do not reliably reflect word grounding. This vector serves as the target for our steering method, described next.
Figure 3: Illustration of TempoControl. During a single denoising step $t$, we extract spatial attention maps $\bar{A}^{t}_{j,i}$ (for word $i$ at temporal index $j$), aggregate to a temporal attention signal $a_{i}^{t}$, and align it with the target mask vector $m_{i}$ via temporal and spatial losses. Gradients $\nabla\mathcal{L}$ are used to update the latent code $z_{t}$.
4 Method
Given a video $v$, our objective is to generate an edited version $v^{\prime}$ such that specific tokens appear according to a temporal condition. To achieve fine-grained temporal control, we employ masks, which provide a more precise and less ambiguous mechanism than textual descriptions. We denote as input a binary mask $m_{i}=[m_{i,1},\ldots,m_{i,T^{\prime}}]$, where each element $m_{i,j}\in\{0,1\}$ indicates whether token $p_{i}$ should appear at temporal position $j$ (see Figure 2, bottom, orange line). In this example, the mask specifies that the token dog should appear only in the second half of the video (i.e., frames 10–20). In general, $m_{i,j}$ may also take continuous values to represent the strength of appearance. In most experiments, we use binary masks, except in the audio-alignment setting, where the mask encodes sound intensity. The main steps of our approach are illustrated in Figure 3 and discussed below. We also provide an algorithmic description in the appendix.
We steer the latent representation during the first $k$ denoising steps by applying $l$ gradient-based updates at each step, optimized with AdamW. Let $z_{t}$ denote the latent at step $t$; the update is:
$z^{\prime}_{t}=z_{t}-\alpha\nabla_{z_{t}}\mathcal{L}^{t},$ (3)
where $\alpha$ is the learning rate. The loss $\mathcal{L}^{t}$ consists of a single main component: a temporal correlation term that aligns the attention pattern with the target mask. While this term is most significant for aligning with the temporal condition, we also include two terms that further improve the results: (ii) a magnitude term that amplifies or suppresses token-level activation, and (iii) an attention entropy regularization term that preserves spatial consistency. We describe each component in detail below.
Temporal correlation term.
The main loss term encourages the temporal attention vector $a_{i}^{t}$ to match the target mask $m_{i}$ using the Pearson correlation:
$\mathcal{L}_{\text{corr}}^{t}=-\frac{\text{Cov}(m_{i},\tilde{a}_{i}^{t})}{\sigma_{m_{i}}\sigma_{\tilde{a}_{i}^{t}}},$ (4)
where $\tilde{a}_{i}^{t}$ is the min–max normalized version of $a_{i}^{t}$, defined as $\tilde{a}_{i}^{t}=\frac{a_{i}^{t}-\min(a_{i}^{t})}{\max(a_{i}^{t})-\min(a_{i}^{t})}$. Here, $\text{Cov}(\cdot,\cdot)$ denotes the empirical covariance, and $\sigma_{m_{i}}$ and $\sigma_{\tilde{a}_{i}^{t}}$ are the corresponding standard deviations. Note that if $m_{i}$ has zero variance, such as when encouraging an object to appear throughout the entire video, the correlation term becomes undefined; in such cases, it is omitted from the loss.
The Pearson correlation measures the linear relationship between the mask $m_{i}$ and the attention term $a_{i}^{t}$, but it does not capture their scale. In our experiments, we observe two phenomena: (i) words that are initially invisible may remain so, and (ii) completely suppressing the appearance of a token is difficult. The following loss term is therefore designed to modulate the temporal strength of attention.
Attention magnitude term.
The magnitude term adjusts the overall strength of attention for each word token $i$ according to the temporal mask $m_{i}$, complementing the correlation-based term by restoring scale sensitivity. We employ the indicator function $\mathbb{1}_{\{\cdot\}}$, which equals one when the specified condition holds and zero otherwise. The positive magnitude component encourages attention in regions where the word is expected to be active:
$\mathcal{L}_{\oplus}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}_{\{m_{i,j}>\tau\}}\cdot a_{i,j}^{t}.$ (5)
The negative magnitude component penalizes attention where the word should not be visible:
$\mathcal{L}_{\ominus}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}_{\{m_{i,j}\leq\tau\}}\cdot a_{i,j}^{t}.$ (6)
The final magnitude loss is defined as the net difference:
$\mathcal{L}_{\text{mag}}^{t}=\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}.$ (7)
Figure 4: Entropy regularization helps preserve semantic fidelity.
Here, $\tau\in[0,1]$ is a threshold that determines whether a given time index is considered active for token $p_{i}$. While $m_{i}$ is often binary (i.e., 0 or 1), it may also take continuous values (e.g., for soft audio-driven conditioning). In such cases, the indicator $\mathbb{1}_{\{m_{i,j}>\tau\}}$ selects time steps where the target activation is sufficiently strong.
Both the correlation and magnitude terms target the temporal objective but can harm spatial consistency. In our experiments, these terms sometimes corrupt object semantics, such as a smartphone turning into a distorted two-screen tablet (Figure 4; more examples in the appendix). To address this and preserve spatial coherence, we introduce an entropy-based regularization term.
Attention entropy regularization.
Optimizing for the temporal objective may compromise spatial properties, often resulting in overly spread attention. We counter this by regularizing the entropy of the spatial attention maps at temporal indices where the word is active (i.e., $m_{i,j}>\tau$), resulting in:
$\mathcal{L}_{\text{entropy}}^{t}=\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot\mathcal{H}(\bar{A}^{t}_{j,i}).$ (8)
Here, $\mathcal{H}(\bar{A})$ denotes the Shannon entropy computed over the spatial dimensions.
In our experiments, we find that discouraging excessive spreading of attention not only improves temporal presence but also enhances overall image quality (see ablations), suggesting that this term may be beneficial for video generation more broadly.
Full objective.
The combined loss at step $t$ is:
$\mathcal{L}^{t}=\mathcal{L}_{\text{corr}}^{t}+\lambda_{1}\mathcal{L}_{\text{magnitude}}^{t}+\lambda_{2}\mathcal{L}_{\text{entropy}}^{t},$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ control the auxiliary terms. The correlation term is sufficient for most cases and provides strong temporal alignment. The magnitude term is helpful when the base model fails to accurately depict a token or when the token must be removed entirely. The entropy term maintains spatial focus and enhances video quality. The objective is optimized over latents during inference, while the model weights remain fixed.
Correlation-based early stopping.
Some videos naturally align more closely with the temporal objective and therefore require fewer optimization steps. To adaptively allocate compute, we introduce a stopping rule based on the Pearson correlation in the temporal loss $\mathcal{L}_{\text{corr}}^{t}$. Specifically, for each denoising step $t$, we perform up to $l$ optimization iterations. If the correlation exceeds a threshold $\tau_{\text{corr}}$, we terminate optimization early for that step; otherwise, we continue refining until the limit is reached.
5 Experiments
Experimental setup.
All experiments are conducted on a single NVIDIA H200 GPU using Wan 2.1-S as the backbone. Due to limited resources, we focus on the smaller model, which already outperforms larger baselines. We optimize the latent representation during the first 10 denoising steps, applying up to 10 gradient updates per step with a learning rate of $5\times 10^{-4}$, increased to $1\times 10^{-3}$ for the two-object setting. We use $\lambda_{1}{=}0.3$ and $\lambda_{2}{=}10$. See the appendix for ablation details.
Baseline.
We evaluate our method on several text-to-video backbones, including Wan 2.1 (1.3B and 14B), Wan 2.2-TI2V (5B), and CogX1.5 (5B). As a baseline, we use each model with explicit temporal cues in the prompt (e.g., “…at the third second …”). The exact phrasing templates are provided in the appendix. Note that we also use the same timing templates in our approach for fair comparison. The lack of dedicated baselines stems from the novelty of our objective, which involves fine-grained temporal control and necessitates the definition of new evaluation metrics, discussed next.
Metrics.
Current benchmarks such as VBench [14] do not measure temporal control. To address this, we introduce Temporal Accuracy, which checks whether each token appears in the correct frames specified by the prompt, with separate scores for presence and absence. VBench relies on the GRiT detector [29], which we find unreliable for small or partially visible objects. Instead, we detect objects using the more recent YOLOv10 [24].
For motion-centric prompts, Temporal Accuracy aligns the prompt-specified action time with the second exhibiting the strongest motion, as determined by per-second optical flow magnitude computed from videos sampled at 4 fps. Full details are provided in the appendix.
We also report results on the VBench Multiple Object benchmark (detecting object appearance throughout the video), and the Imaging Quality metric that uses a perceptual image-quality model.
Benchmark.
In the single-object setup, we utilize 80 YOLOv10 classes, which span animals, everyday objects, vehicles, and more. For the two-object setting, we use 82 object pairs based on the VBench multiple-object benchmark. For movement, we curate 100 types of actions. Temporal masks are generated together with the prompt templates to ensure alignment. Full prompt patterns and categories are provided in the appendix.
Table 1: Comparison across one-object, two-object, and movement temporal-control tasks. Baselines use text-only temporal cues, while Ours applies TempoControl on Wan2.1 (1.3B). Bold values indicate the best result in each block.
Method # Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Imaging Quality
One Object
CogX1.5 5B 61.43% 48.64% 75.50% 65.36%
Wan2.1-S 1.3B 63.94% 67.38% 60.50% 53.76%
Wan2.1-L 14B 64.23% 55.34% 74.00% 66.89%
Wan2.2 5B 78.02% 87.42% 68.00% 48.13%
Ours 1.3B 82.50% 83.25% 81.75% 56.51%
Two Objects
CogX1.5 5B 24.68% 46.46% 4.88% 63.78%
Wan2.1-S 1.3B 37.50% 45.85% 29.15% 68.56%
Wan2.1-L 14B 40.07% 54.15% 27.27% 69.01%
Wan2.2 5B 47.05% 56.86% 36.59% 67.83%
Ours 1.3B 52.93% 58.05% 47.80% 71.09%
Movement
CogX1.5 5B 21% 61.39%
Wan2.1-S 1.3B 19% 60.46%
Wan2.1-L 14B 18% 61.31%
Wan2.2 5B 23% 60.61%
Ours 1.3B 57% 63.38%
5.1 Quantitative Results
Single object temporal control.
In Table 1 (One Object), our method improves temporal accuracy by 18.56%, with gains of 15.87% in absence accuracy and 21.25% in presence accuracy over the Wan2.1-S baseline. Imaging Quality also increases by about 3%.
Larger baseline models achieve higher imaging quality, as expected, yet all fail to provide reliable temporal control from text alone. Wan2.2’s strong absence accuracy largely reflects degraded image quality, which causes detectors to miss objects. In a separate experiment, we tested removing the timing cues, which increased its Imaging Quality to 57.78%, indicating that it is not robust to temporal instructions; details are in the appendix.
Two objects temporal control.
In Table 1 (Two Objects), the two-object setting is notably challenging. Wan2.1-S baseline achieves only 37.50% temporal accuracy, while our method improves this to 52.93% (+15.43). Absence and presence accuracies also increase by 12.20% and 18.65%, respectively, and Imaging Quality improves by 2.53%.
Larger models also show similar degradation in temporal control in the two-object setting. Conversely, Imaging Quality is consistently higher across all methods, suggesting a possible bias in the perceptual metric toward multi-object scenes. Our method achieves the highest Imaging Quality.
Movement temporal control.
Table 1 (Movement) evaluates temporal accuracy when conditioning is indicative of a peak of action. The temporal accuracy measures the timing of when the action peaks (see Metrics). Wan2.1-S baseline achieves only 19% temporal accuracy, indicating that the model does not follow action timing from text alone. TempoControl improves this to 57%, showing a substantial gain in aligning motion onset with the prompt.
Low performance in movement conditioning is consistent across all models. Our Imaging Quality remains the best in this setup as well.
Multiple objects benchmark.
We also report results on the VBench Multiple Objects benchmark in Table 2, which tests the ability to generate two objects described in the input text throughout the entire video. We observe improved performance on the Multiple Object metric (76.34% vs. 74.13%). Detailed results are provided in the appendix.
Ablation.
Table 3 shows that each component of our objective contributes differently to temporal control and visual quality. Pearson correlation (C) provides the strongest temporal alignment, driven largely by improvements in T. Absence Accuracy, but it also reduces Imaging Quality. Closer inspection suggests that degraded visuals can cause detectors to miss objects, artificially inflating absence scores. Entropy regularization (E) improves presence accuracy and stabilizes spatial attention. Combining both (C+E) yields balanced gains across metrics. Notably, using entropy alone produces the highest Imaging Quality (59.52%), outperforming even the text-only baseline, suggesting that entropy-based attention smoothing is broadly beneficial for video generation. Additional ablations are provided in the appendix.
Human evaluation.
We conducted a user study with 50 computer science graduate students, each comparing 16 video pairs (8 one-object and 8 two-object) generated by TempoControl and Wan 2.1 with text-based temporal cues. Annotators judged temporal accuracy and visual quality, or indicated no difference, which we asked them to use sparingly (full protocol details are provided in the appendix). As shown in Table 4, our method is preferred for temporal accuracy (61.51% vs. 16.94%) and also improves visual quality. Notably, 21.55% of temporal comparisons reported no preference, reflecting that humans struggle to perceive minor timing differences.
Table 2: Full results for the multiple-object benchmark. Best results per column are in bold.
Method Multiple Object (GRiT) Multiple Object (YOLO) Imaging Quality
Text 74.13% 61.54% 70.25%
Ours 76.37% 65.73% 70.21%
Table 3: Ablation results for the one-object setting. C = Pearson correlation, E = Entropy.
Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Imaging Quality
Text 63.94% 67.38% 60.50% 53.76%
Only C 81.19% 91.50% 70.88% 50.96%
Only E 72.94% 66.88% 79.00% 59.52%
C + E 78.38% 77.38% 79.38% 57.60%
Ours 82.50% 83.25% 81.75% 56.51%
Table 4: User study comparing our method with Wan2.1 in terms of temporal accuracy and visual quality.
Question Type Ours Baseline Same
Temporal Accuracy 61.51% 16.94% 21.55%
Visual Quality 62.66% 25.99% 11.35%
Figure 5: Examples generated by Wan 2.1 (left) and our temporally conditioned method (right) for prompts with timing constraints. Prompts are simplified for brevity. Shown are frames from the first, middle, and final thirds of each video.
Figure 6: Examples of video alignment to an audio signal.
5.2 Qualitative Results
In Figure 5, each row illustrates a different temporal-control task. The first row shows a one-object example: for the prompt “an apple appears at the fourth second,” Wan 2.1 places the apple near the middle of the five-second video, whereas our method correctly places it near the end. The second row shows a two-object example: for “a car and a bicycle, car in the second half,” Wan 2.1 displays the car at the beginning instead of the specified time. The third and fourth rows depict movement-based prompts. For a magician pulling a rabbit at the last second, Wan 2.1 shows the rabbit throughout the video, whereas our method reveals it only at the end. For a tiger leaping in the first second, Wan 2.1 fails to time the leap, whereas our method accurately aligns the motion with the prompt. Additional examples are included in the supplementary material.
5.3 Audio-Visual Alignment
We also explore zero-shot alignment of video to external audio by using a preprocessed audio envelope as the temporal condition. Given a waveform $y$, we compute its onset-strength envelope $s_{t}$, pool it to the number of frames, and normalize it to $[0,1]$. Peaks are preserved through thresholding and otherwise smoothed with a Gaussian filter:
$\tilde{s}_{t}=\begin{cases}\hat{s}_{t},&\hat{s}_{t}\geq\tau_{audio},\\ \text{GaussianSmooth}(\hat{s}_{t};\sigma),&\text{otherwise}.\end{cases}$ (10)
Figure 6 shows examples where video motion aligns with peaks in the audio envelope, with additional results in the appendix. While tested on simple cases with clear peaks, these results suggest that audio-driven temporal conditioning is a promising approach for achieving audio-video alignment without requiring training on paired datasets.
Limitations.
TempoControl has two main limitations. First, it increases inference time to approximately 460 seconds per video, compared to 170 seconds for vanilla generation. Nonetheless, it remains an inference-only approach that requires no additional supervision, thereby avoiding the degradation in robustness and diversity often seen in supervised training. Second, our conditioning can introduce mild perceptual shifts in target attributes (for example, altering the color of a cat; see Figure 1). While consistency is often maintained, our current objectives do not fully guarantee it, making this an interesting direction for future work.
Temporal control in text-to-video models.
We observe that adding explicit timing cues to current text-to-video models does not improve temporal accuracy and often degrades visual quality. For example, in Wan 2.1 (one-object setting), temporal accuracy remains at 64% while Imaging Quality drops from 56% to 53%. This highlights a broader limitation: existing models lack reliable temporal grounding from text alone. For fairness, we keep the same prompt format for all baselines, though omitting explicit timing slightly improves our Imaging Quality (56% vs. 53%) without affecting Temporal Accuracy (details in the appendix).
5.4 Conclusions
Current text-to-video models generate semantically accurate and high-quality videos, but they lack precise control over the temporal appearance of objects and actions. Our results show that existing text-conditioning approaches are insufficient for temporal control. To address this gap, we present a data-free temporal conditioning technique, which maintains video quality while enabling fine-grained temporal control through attention steering. We hope our approach will inspire new solutions to tasks such as audio-video alignment, as discussed in this work.
Supplementary Material for TempoControl: Temporal Attention Guidance for Text-to-Video Models
This supplementary material includes algorithmic pseudocode, ablation studies, experimental design details, application-specific configurations, extended quantitative and qualitative results, and the human-evaluation setup. Videos and code are also provided in the supplementary package.
The sections of the supplementary material are as follows:
  1. 1.
    Algorithm pseudocode (Section .5).
  2. 2.
    Experimental setup and hyperparameters (Section .6).
  3. 3.
    Configuration details for each application setting (Section .7).
  4. 4.
    Human evaluation protocol and interface (Section .8).
  5. 5.
    Extended quantitative results for multi-object grounding (Section .9).
  6. 6.
    Ablation study of loss components and qualitative failure modes (Section .10).
  7. 7.
    Analysis of prompt phrasing and its impact on performance (Section .11).
  8. 8.
    Additional qualitative results with sampled frames (Section .12).
.5 Algorithm Pseudo-code
Algorithm 1 outlines our inference-time optimization procedure. Starting from noise $z_{T}$, we iteratively steer the latent code during the first $k$ denoising steps of the diffusion process. At each step $t$, we compute cross-attention maps from the video-text model $\operatorname{DiT}$ and derive the scalar temporal attention $a_{i}^{t}$ for each token $p_{i}$.
The optimization loss $\mathcal{L}$ combines three components: correlation alignment $\mathcal{L}_{\text{c}}^{t}$, magnitude modulation $\mathcal{L}_{\text{enr}}^{t}$, entropy regularization and $\mathcal{L}_{\text{ent}}^{t}$. The latent code is updated via gradient descent for up to $l$ steps per denoising timestep, or until the Pearson correlation exceeds a threshold $\tau_{\text{corr}}$. After optimization, the standard denoising process resumes to produce the final latent $z_{0}$.
Algorithm 1 TempoControl: Temporal Attention-Guided Inference-Time Steering
Input: Prompt $\mathcal{P}$, tokens $p_{i}$ with temporal masks $m_{i}$, total diffusion steps $T$, optimized steps $k$, max updates per step $l$, correlation threshold $\tau_{\text{corr}}$, model $\operatorname{DiT}$
Output: Final latent $z_{0}$
1:Initialize $z_{T}\sim\mathcal{N}(0,I)$
2:for $t=T,T{-}1,\dots,T{-}k{+}1$ do
3:  for $u=1$ to $l$ do
4:   Extract cross-attentions $A^{t}$ from $\operatorname{DiT}(z_{t},\mathcal{P},t)$
5:   Average heads and layers: $\bar{A}^{t}$
6:   $\mathcal{L}\leftarrow 0$
7:   for each $p_{i}$ with $m_{i}$ do
8:     $a_{i}^{t}\leftarrow[\langle\bar{A}^{t}_{j,i}\rangle_{x,y}]_{j=1}^{T^{\prime}}$
9:     $\mathcal{L}_{\text{c}}^{t}\leftarrow-\text{Corr}(m_{i},\text{minmax}(a_{i}^{t}))$
10:     $\mathcal{L}_{\oplus}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot a_{i,j}^{t}$
11:     $\mathcal{L}_{\ominus}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}\leq\tau]\cdot a_{i,j}^{t}$
12:     $\mathcal{L}_{\text{mag}}^{t}\leftarrow\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}$
13:     $\mathcal{L}_{\text{ent}}^{t}\leftarrow\frac{1}{T^{\prime}}\sum_{j=1}^{T^{\prime}}\mathbb{1}[m_{i,j}>\tau]\cdot\mathcal{H}(\bar{A}^{t}_{j,i})$
14:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\text{c}}^{t}+\lambda_{1}\mathcal{L}_{\text{enr}}^{t}+\lambda_{2}\mathcal{L}_{\text{ent}}^{t}$
15:   end for
16:   $z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\nabla_{z_{t}}(\mathcal{L}/N)$
17:   if All $\mathcal{L}_{\text{corr}}^{t}\leq\tau_{\text{corr}}$ then
18:     break
19:   end if
20:   $z_{t}\leftarrow z_{t}^{\prime}$
21:  end for
22:  $z_{t-1}\leftarrow\operatorname{DiT}(z_{t}^{\prime},\mathcal{P},t)$
23:end for
24:Continue denoising using the standard denoising diffusion process for $t=T{-}k,\dots,1$
25:return $z_{0}$
.6 Experimental Setup
Hyperparameters were tuned early on using a small validation split tailored to each task. Although large-scale ablations were constrained by computational cost, the method proved robust across settings, requiring only minor adjustments.
All experiments use Wan 2.1 as the backbone model. We apply a classifier-free guidance scale of 6 and a sample shift of 3 during generation.
For single-object scenes, we use $\lambda_{1}{=}0.3$, $\lambda_{2}{=}10$, a learning rate of $5{\times}10^{-4}$, a Pearson stopping threshold of $\tau_{\text{corr}}{=}0.9$, and perform optimization over the first l=10l{=}10 denoising steps with up to k=10k{=}10 gradient updates per step.
For two-object scenes, which demand stronger guidance, we increase the learning rate to $10^{-3}$ to improve convergence. For motion-centric prompts, we relax the Pearson threshold to $\tau_{\text{corr}}{=}0.85$ to enable more flexible optimization.
.7 Applications Setup
Single-Object Temporal Control.
To evaluate temporal grounding for individual objects, we construct a dataset based on the 80 object categories detectable by the YOLO detector. These categories cover common entities such as animals, vehicles, household items, and everyday objects. The complete class list is provided below:
person, bicycle, car, motorcycle, airplane, bus, train, truck, boat, traffic light, fire hydrant, stop sign, parking meter, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, chair, couch, potted plant, bed, dining table, toilet, tv, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, sink, refrigerator, book, clock, vase, scissors, teddy bear, hair drier, toothbrush.
Videos are generated from prompts of the form ‘An empty scene. Suddenly, $\langle o_{1}\rangle$ appears out of nowhere, drawing all attention.’ Object classes are uniformly sampled across different temporal intervals.
Two-Object Temporal Control.
We evaluate TempoControl on 82 prompts derived from the VBench protocol, each consisting of a pair of objects. The complete numbered list of pairs is provided below:
  1. 1.
    a bird and a cat
  2. 2.
    a cat and a dog
  3. 3.
    a dog and a horse
  4. 4.
    a horse and a sheep
  5. 5.
    a sheep and a cow
  6. 6.
    a cow and an elephant
  7. 7.
    an elephant and a bear
  8. 8.
    a bear and a zebra
  9. 9.
    a zebra and a giraffe
  10. 10.
    a giraffe and a bird
  11. 11.
    a chair and a couch
  12. 12.
    a couch and a potted plant
  13. 13.
    a potted plant and a tv
  14. 14.
    a tv and a laptop
  15. 15.
    a laptop and a remote
  16. 16.
    a remote and a keyboard
  17. 17.
    a keyboard and a cell phone
  18. 18.
    a cell phone and a book
  19. 19.
    a book and a clock
  20. 20.
    a clock and a backpack
  21. 21.
    a backpack and an umbrella
  22. 22.
    an umbrella and a handbag
  23. 23.
    a handbag and a tie
  24. 24.
    a tie and a suitcase
  25. 25.
    a suitcase and a vase
  26. 26.
    a vase and scissors
  27. 27.
    scissors and a teddy bear
  28. 28.
    a teddy bear and a frisbee
  29. 29.
    a frisbee and skis
  30. 30.
    skis and a snowboard
  31. 31.
    a snowboard and a sports ball
  32. 32.
    a sports ball and a kite
  33. 33.
    a kite and a baseball bat
  34. 34.
    a baseball bat and a baseball glove
  35. 35.
    a baseball glove and a skateboard
  36. 36.
    a skateboard and a surfboard
  37. 37.
    a surfboard and a tennis racket
  38. 38.
    a tennis racket and a bottle
  39. 39.
    a bottle and a chair
  40. 40.
    an airplane and a train
  41. 41.
    a train and a boat
  42. 42.
    a boat and an airplane
  43. 43.
    a bicycle and a car
  44. 44.
    a car and a motorcycle
  45. 45.
    a motorcycle and a bus
  46. 46.
    a bus and a traffic light
  47. 47.
    a traffic light and a fire hydrant
  48. 48.
    a fire hydrant and a stop sign
  49. 49.
    a stop sign and a parking meter
  50. 50.
    a parking meter and a truck
  51. 51.
    a truck and a bicycle
  52. 52.
    a toilet and a hair drier
  53. 53.
    a hair drier and a toothbrush
  54. 54.
    a toothbrush and a sink
  55. 55.
    a sink and a toilet
  56. 56.
    a wine glass and a chair
  57. 57.
    a cup and a couch
  58. 58.
    a fork and a potted plant
  59. 59.
    a knife and a tv
  60. 60.
    a spoon and a laptop
  61. 61.
    a bowl and a remote
  62. 62.
    a banana and a keyboard
  63. 63.
    an apple and a cell phone
  64. 64.
    a sandwich and a book
  65. 65.
    an orange and a clock
  66. 66.
    broccoli and a backpack
  67. 67.
    a carrot and an umbrella
  68. 68.
    a hot dog and a handbag
  69. 69.
    a pizza and a tie
  70. 70.
    a donut and a suitcase
  71. 71.
    a cake and a vase
  72. 72.
    an oven and scissors
  73. 73.
    a toaster and a teddy bear
  74. 74.
    a microwave and a frisbee
  75. 75.
    a refrigerator and skis
  76. 76.
    a bicycle and an airplane
  77. 77.
    a car and a train
  78. 78.
    a motorcycle and a boat
  79. 79.
    a person and a toilet
  80. 80.
    a person and a hair drier
  81. 81.
    a person and a toothbrush
  82. 82.
    a person and a sink
Prompts follow the structure: ‘The video begins with a serene view centered on $\langle o_{1}\rangle$, with no sign of $\langle o_{2}\rangle$. In the second half, $\langle o_{2}\rangle$ unexpectedly appears, altering the dynamic of the scene.’
A static control signal (all ones) is used for $\langle o_{1}\rangle$. We sample 20 evenly spaced frames and require $\langle o_{1}\rangle$ to be present throughout. A frame is counted as successful if both objects are detected when $\langle o_{2}\rangle$ is active, and only $\langle o_{1}\rangle$ is detected when $\langle o_{2}\rangle$ is inactive.
Motion Temporal Control.
To evaluate fine-grained temporal modulation of actions, we generate 100 videos using prompts of the form: ‘A video of $\langle s\rangle$ $\langle v\rangle$ $\langle a\rangle$ with a strong movement at the $\langle t\rangle$ second,’ where $\langle s\rangle$, $\langle v\rangle$, $\langle a\rangle$, and $\langle t\rangle$ denote the subject, verb, adverb, and time reference, respectively.
The full list of subject–verb–adverb triplets is given below:
  1. 1.
    a cat pouncing silently
  2. 2.
    a dancer twirling elegantly
  3. 3.
    a bird flapping frantically
  4. 4.
    a lion roaring loudly
  5. 5.
    a wave crashing violently
  6. 6.
    a ballerina leaping lightly
  7. 7.
    a helicopter spinning noisily
  8. 8.
    a snake slithering smoothly
  9. 9.
    a bear growling deeply
  10. 10.
    a man running fast
  11. 11.
    a woman laughing brightly
  12. 12.
    a child crawling slowly
  13. 13.
    a monkey swinging playfully
  14. 14.
    a firework exploding suddenly
  15. 15.
    a wolf howling hauntingly
  16. 16.
    a dolphin diving swiftly
  17. 17.
    a plane ascending steadily
  18. 18.
    a kangaroo hopping energetically
  19. 19.
    a fox darting quickly
  20. 20.
    a shadow moving mysteriously
  21. 21.
    a squirrel scampering nervously
  22. 22.
    a truck crashing loudly
  23. 23.
    a child screaming unexpectedly
  24. 24.
    a penguin waddling adorably
  25. 25.
    a frog croaking rhythmically
  26. 26.
    a leaf drifting slowly
  27. 27.
    a shark circling menacingly
  28. 28.
    a comet streaking brightly
  29. 29.
    a dancer collapsing dramatically
  30. 30.
    a boxer punching fiercely
  31. 31.
    a tree swaying gently
  32. 32.
    a man lifting weights powerfully
  33. 33.
    a windmill rotating steadily
  34. 34.
    a girl skipping cheerfully
  35. 35.
    a car drifting dangerously
  36. 36.
    a train accelerating fast
  37. 37.
    a snake striking quickly
  38. 38.
    a fire burning intensely
  39. 39.
    a glacier cracking slowly
  40. 40.
    a bee buzzing constantly
  41. 41.
    a deer sprinting fearfully
  42. 42.
    a volcano erupting violently
  43. 43.
    a runner collapsing from exhaustion
  44. 44.
    a gymnast flipping smoothly
  45. 45.
    a rocket launching thunderously
  46. 46.
    a drummer hitting rapidly
  47. 47.
    a magician vanishing mysteriously
  48. 48.
    a spider crawling delicately
  49. 49.
    a tiger pacing restlessly
  50. 50.
    a surfer balancing skillfully
  51. 51.
    a leopard growling softly
  52. 52.
    a swimmer diving gracefully
  53. 53.
    a baby clapping happily
  54. 54.
    a mime gesturing expressively
  55. 55.
    a crow cawing sharply
  56. 56.
    a goat headbutting suddenly
  57. 57.
    a girl blowing bubbles gently
  58. 58.
    a chef chopping rapidly
  59. 59.
    a horse shaking its mane proudly
  60. 60.
    a robot malfunctioning erratically
  61. 61.
    a meteor falling fast
  62. 62.
    a snake coiling tightly
  63. 63.
    a rabbit thumping nervously
  64. 64.
    a man collapsing dramatically
  65. 65.
    a car braking suddenly
  66. 66.
    a hawk diving precisely
  67. 67.
    a wolf snarling viciously
  68. 68.
    a plane landing smoothly
  69. 69.
    a firetruck speeding urgently
  70. 70.
    a chimpanzee clapping playfully
  71. 71.
    a dancer stomping rhythmically
  72. 72.
    a snail inching slowly
  73. 73.
    a lioness crouching quietly
  74. 74.
    a storm cloud rolling ominously
  75. 75.
    a bee landing precisely
  76. 76.
    a magician pulling a rabbit suddenly
  77. 77.
    a glacier melting steadily
  78. 78.
    a jellyfish floating gracefully
  79. 79.
    a tree falling loudly
  80. 80.
    a drummer drumming fiercely
  81. 81.
    a torch blazing brightly
  82. 82.
    a woman spinning dramatically
  83. 83.
    a child running barefoot
  84. 84.
    a fox sniffing cautiously
  85. 85.
    a man throwing a ball forcefully
  86. 86.
    a bear shaking off water heavily
  87. 87.
    a lion walking regally
  88. 88.
    a horse rearing suddenly
  89. 89.
    a falcon gliding silently
  90. 90.
    an elephant raising its trunk highly
  91. 91.
    a dog barking forcefully
  92. 92.
    a man sneezing powerfully
  93. 93.
    a child jumping joyfully
  94. 94.
    a tiger leaping fiercely
  95. 95.
    a woman spinning gracefully
  96. 96.
    a robot marching stiffly
  97. 97.
    a horse galloping wildly
  98. 98.
    an eagle soaring majestically
  99. 99.
    a flame flickering rapidly
  100. 100.
    a skateboarder flipping expertly
The verb is modulated using a binary temporal mask (set to 1 only at time $\langle t\rangle$), while the subject remains active throughout the video.
.8 Human Evaluation Protocol
To assess temporal alignment and visual quality, we conducted a human preference study using side-by-side video comparisons. All annotators were graduate students conducting research in computer vision or closely related areas.
For each prompt, two videos were shown: one generated by our method and one by the text-only baseline, with their order (Video A or Video B) randomized to avoid bias.
Annotators were instructed to read the prompt and answer two questions: (1) Which video is more temporally accurate (i.e., does the object appear at the correct time and remain visible)? (2) Which video is more visually appealing? A “Same (use sparingly)” option was available when no meaningful difference was perceived.
After the experiments, several annotators noted that judging temporal control could be challenging when the difference in timing was only 1–2 seconds. They reported that knowing the exact expected trigger time from the control mask would have helped them assess temporal accuracy more reliably. While we chose not to repeat the study, we include this observation as guidance for future temporal evaluation protocols.
Example screenshots of the evaluation interface are shown in Figures 7 and 8.
Figure 7: Evaluation interface showing the prompt and two videos (A and B) for comparison.
Figure 8: Second example of the human evaluation interface.
.9 Multi-Object Evaluation
We report extended results on multi-object text-to-video generation using the VBench benchmark. In this setting, prompts reference multiple entities (e.g., ‘A dog and a cat’), and the evaluation measures whether all specified objects appear consistently throughout the video. Each video is sampled at 16 evenly spaced frames, and a frame is counted as successful if both target objects are detected; the final score is the average per-frame success rate. In the context of our method, this setup corresponds to applying a constant mask of one for both objects across all frames.
For completeness, we briefly summarize the VBench metrics reported in Table 5:
  • Multiple-Objects Accuracy (GriT / YOLO). Measures whether all referenced objects appear correctly in each sampled frame, using open-vocabulary detectors (GriT) or standard detectors (YOLO). Higher values indicate more reliable object visibility and grounding.
  • Subject Consistency. Quantifies how consistently the appearance of the main subject is preserved across frames. It is computed using feature similarity both between consecutive frames and relative to the first frame of the video; higher values indicate a more stable and coherent subject identity.
  • Background Consistency. Measures the stability of the background across the video. High scores indicate coherent scene structure without abrupt changes.
  • Dynamic Degree. Captures the amount of motion or temporal change in the video. The metric is binary at the video level: a video is marked as dynamic if its average optical flow surpasses a predefined threshold. Higher values indicate more motion, while lower values reflect greater temporal stability.
Table 5 summarizes the results. Our method achieves substantial gains in object-centric metrics: multi-object accuracy improves from 74.1% to 76.4% (GriT) and from 61.5% to 65.7% (YOLO), demonstrating better grounding and more reliable visibility of both entities. We also observe improvements across consistency metrics, including subject consistency (from 97.2% to 97.8%) and background consistency (from 97.6% to 98.1%).
The Dynamic Degree decreases from 30.5% to 18.8%, reflecting a trade-off: our method promotes temporal stability, often leading to more consistent object appearance across the video, at the cost of reduced motion. This highlights a limitation of this evaluation protocol: VBench counts an object as correct only when it is present in every sampled frame; hence, the benchmark implicitly favors unnatural videos in which objects persist throughout the sequence.
Table 5: Full results for Multiple object benchmark results. Best results per column are in bold.
Method Multiple Object (GriT) Multiple Object (YOLO) Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Text 74.13% 61.54% 97.22% 97.56% 99.18% 30.49% 62.84% 70.25% Ours 76.37% 65.73% 97.81% 98.10% 99.40% 18.78% 62.52% 70.21%
.10 Ablations
We conduct an ablation study to analyze the influence of the magnitude weight $\lambda_{1}$ and the entropy weight $\lambda_{2}$ on temporal control and visual quality. The results reveal clear and complementary effects of the two loss components. Increasing the magnitude term $\lambda_{1}$ reduces Imaging Quality while improving T. Absence Accuracy. This suggests that stronger magnitude guidance more strictly enforces the temporal mask, but may also reduce overall visual fidelity.
In contrast, increasing the entropy term $\lambda_{2}$ improves Imaging Quality and raises T. Presence Accuracy, but it decreases T. Absence Accuracy, indicating that stronger entropy regularization encourages more persistent object appearance.
Overall, our chosen hyperparameters ($\lambda_{1}=0.3$, $\lambda_{2}=10$) provide a balanced tradeoff between these competing effects, achieving strong Temporal Accuracy while maintaining high Imaging Quality.
Table 6: Ablation results for the one object setting. All variants use the same temporal Pearson term; we ablate the magnitude weight $\lambda_{1}$ and the entropy weight $\lambda_{2}$. Best scores per column are in bold.
(a) Effect of magnitude weight $\lambda_{1}$ with fixed entropy weight $\lambda_{2}=10$.
Method
Temporal
Accuracy
T. Absence
Accuracy
T. Presence
Accuracy
Imaging
Quality
Text baseline 63.94% 67.38% 60.50% 53.76%
Ours ($\lambda_{1}{=}0.3,\lambda_{2}{=}10$) 82.50% 83.25% 81.75% 56.51%
$\lambda_{1}{=}0.1,\lambda_{2}{=}10$ 79.50% 80.62% 78.38% 56.37%
$\lambda_{1}{=}0.65,\lambda_{2}{=}10$ 81.06% 86.50% 75.62% 53.94%
$\lambda_{1}{=}1.0,\lambda_{2}{=}10$ 83.12% 89.00% 77.25% 53.74%
(b) Effect of entropy weight $\lambda_{2}$ with fixed magnitude weight $\lambda_{1}=0.3$.
Method
Temporal
Accuracy
T. Absence
Accuracy
T. Presence
Accuracy
Imaging
Quality
Text baseline 63.94% 67.38% 60.50% 53.76%
Ours ($\lambda_{1}{=}0.3,\lambda_{2}{=}10$) 82.50% 83.25% 81.75% 56.51%
$\lambda_{1}{=}0.3,\lambda_{2}{=}1$ 80.38% 91.25% 69.50% 52.01%
$\lambda_{1}{=}0.3,\lambda_{2}{=}5$ 81.88% 88.50% 75.25% 54.04%
$\lambda_{1}{=}0.3,\lambda_{2}{=}15$ 81.56% 81.62% 81.50% 56.31%
Figure 9 illustrates ablation results for the entropy term, revealing the limitation of using the Pearson correlation term alone without entropy regularization. While Pearson-based optimization successfully forces the object to appear at the desired time, it often alters the object’s semantics, leading to noticeable inconsistencies. For example, in the “Cake” row, the Pearson-only variant transforms the cake into a noticeably different shape and texture; in the “Sandwich” row, the sandwich becomes distorted across frames; and in the “Book” row, the open book becomes a closed stack with altered proportions. Similar semantic drift is observed in the “Keyboard,” “Oven,” and “Stop sign” examples.
These results highlight that Pearson correlation alone can enforce temporal timing but does not sufficiently constrain the appearance space, resulting in unwanted semantic changes. The entropy regularization helps mitigate those artifacts.
Figure 9: Comparison between Wan2.1 and Pearson-only optimization for single-object temporal control. Pearson correlation alone often satisfies the temporal constraint but alters object semantics, producing distorted or inconsistent appearances (e.g., the cake becoming misshaped, the sandwich warping, the book changing structure, or the keyboard turning into a laptop). Adding entropy regularization prevents these failures by stabilizing the object’s identity across time.
.11 Limitations of Explicit Temporal Cues in State-of-the-Art Text-to-Video Models
We analyze the effect of using explicit temporal phrasing (e.g., ‘in the third second’, ‘in the second part of the video’) in the prompt for the one-object setup. Table 7 compares Wan 2.1, Wan 2.2, and our method, each evaluated with and without such phrasing.
For Wan 2.1, explicit temporal cues do not improve temporal grounding. In fact, removing them increases Temporal Accuracy (63.94% to 65.18%) and leads to a clear gain in Imaging Quality (53.76% to 59.99%), indicating that the model struggles to interpret precise timing expressions and performs better without them.
Wan 2.2 shows a more nuanced pattern: it handles temporal phrasing better than Wan 2.1, likely due to improved data or training techniques. However, the drop in quality when explicit timing is introduced is still significant. Imaging Quality decreases sharply from 57.78% (without time) to 48.13% (with time), showing that temporal phrasing introduces confusion even for stronger models. Presence Accuracy also falls (79.50% to 68%), further highlighting that explicit timing signals continue to degrade performance.
Our method remains the most robust and accurate across settings. Across all conditions, our method provides the most reliable and precise temporal control, regardless of prompt style. Interestingly, although our approach builds on Wan 2.1 as the base model, explicit temporal phrasing does increase the temporal accuracy. This suggests that our optimization procedure may curate the base model’s timing capabilities.
Table 7: Effect of temporal phrasing in prompts for the one-object setup. Removing explicit timing improves video quality.
Method Temp. Acc. T. Abs. Acc. T. Pres. Acc. Imaging Quality
Wan2.1 with time 63.94% 67.38% 60.50% 53.76%
Wan2.1 without time 65.18% 65.12% 65.25% 59.99%
Wan2.2 with time 78.02% 87.42% 68.00% 48.13%
Wan2.2 without time 75.24% 71.25% 79.50% 57.78%
Ours with time 82.50% 83.25% 81.75% 56.51%
Ours without time 78.75% 82.25% 75.25% 57.14%
.12 Additional Qualitative Examples
We provide qualitative results across the one-object, two-object, and motion-based setups. For each example, we show sampled frames from the generated videos, illustrating the temporal grounding of object appearances. Prompts are abbreviated for brevity, indicating the key object and its expected time of appearance. Note that once introduced, the object is expected to persist until the end of the video. Videos can be viewed on the project page.
a baseball glove and a skateboard
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a bicycle and a car
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a bicycle and an airplane
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a bird and a cat
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a boat and an airplane
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a book and a clock
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a bowl and a remote
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a bus and a traffic light
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a cake and a vase
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a car and a motorcycle
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a cat and a dog
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a cell phone and a book
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a donut and a suitcase
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Two Objects Text vs Ours.
TextOurs
a giraffe and a bird
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a handbag and a tie
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a horse and a sheep
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a laptop and a remote
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a motorcycle and a boat
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a motorcycle and a bus
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a person and a toothbrush
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a pizza and a tie
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a potted plant and a tv
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a spoon and a laptop
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a tennis racket and a bottle
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a toaster and a teddy bear
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a truck and a bicycle
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
a zebra and a giraffe
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 11: Two Objects Text vs Ours.
TextOurs
airplane 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
apple 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
backpack 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
bench 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
bicycle 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
bowl 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
bus 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
cell phone last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
clock last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
cow 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
dining table last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
dog last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
donut 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
fork 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 12: One Object Text vs Ours.
TextOurs
hair drier last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
handbag 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
kite 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
knife 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
scissors last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
sheep 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
skateboard 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
snowboard 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
sports ball 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
suitacase 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
toilet last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
truck 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
umbrella 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 13: One Object Text vs Ours.
TextOurs
bee buzzing 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
bird flapping 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
child crawling 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
child jumping 1st second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
child screaming 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
chimpanze clapping 1st second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
crow cawing 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
falcon gliding last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
fox darting 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
horse shaking its mane 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
jellyfish floating last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
kangaroo hopping 2nd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
leopard growling 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 14: Movement Text vs Ours.
TextOurs
lion roaring 1st second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
lion walking last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
magician pulling rabbit last second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
penguin waddling 1st second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
rabbit thumping 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
robot marching 1st second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
rocket launching 3rd second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
surfer balancing 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
swimmer diving 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
wolf snarling 4th second
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 15: Movement Text vs Ours.
TextOurs
elephant ex1
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
elephant ex2
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
lightning ex1
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
lightning ex2
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
rooster ex1
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
rooster ex2
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 16: Audio-video Alignment Text vs Ours.